OpenStudy (anonymous):

One factor of \[P(x)=4x^3+8x^2-ax-b\] is x+2. When the polynomial is divided by x-1, the remainder is -15. Find a and b, and hence factor the polynomial completely.

OpenStudy (anonymous):

I think I would have been okay if it was one term... but two terms is confusing me on what I should do first...

OpenStudy (anonymous):

@Limitless Do you know how to answer this question?

OpenStudy (anonymous):

Ok.. So, do you know synthetic division?

OpenStudy (anonymous):

Yes... but I don't know where to use it...

OpenStudy (anonymous):

I think this will take me a moment. Just a sec.

OpenStudy (anonymous):

Okay :).

OpenStudy (anonymous):

@limitless did you get stuck? @satellite73 Do you know how to do this?

OpenStudy (anonymous):

I am quite confused by this problem. I may not arrive at a solution.

OpenStudy (anonymous):

Sigh... and I am being tested on this in the morning... :'(

OpenStudy (anonymous):

Disregard that, I (am so tempted to use a meme here) was overthinking. \[\frac{P(x)}{x+2}=4x^2-\frac{ax+b}{x+2}\] Since \(x+2\) is a factor, it implies \(a=1\) and \(b=2\). Does this make sense?

OpenStudy (anonymous):

I apologize! It should be expressed \[\frac{P(x)}{x+2}=4x^2+\frac{-ax-b}{x+2},\] which implies \(a=-1\) and \(b=-2.\)

OpenStudy (anonymous):

Not at all... lol

OpenStudy (anonymous):

:\...why did you express it that way?

OpenStudy (anonymous):

\(p(-2)=0\) and \(p(1)=-15\) you get two equations in two unknowns

OpenStudy (anonymous):

Okay, here is why that logic works. When we say something is a factor of a polynomial, it must divide in a way such that there is no rational function. So we can't have the expression \[\frac{-ax-b}{x+2}\] in our division because this is a rational function. So, we must find the \(a\) and \(b\) such that \(-ax-b=x+2\). This is how I got that \(a=-1\) and (b=-2\).

OpenStudy (anonymous):

although i assume @Limitless is correct, i would solve two equations

OpenStudy (anonymous):

@satellite73 how does that allow me to solve for a and b? and @Limitless I get confused when math is in two much words...

OpenStudy (anonymous):

replace \(x\) by 1, set result equal to \(-15\) you get \[4+8-a-b=-15\] \[12-a-b=-15\] \[-a-b=-27\] or \[a+b=27\]as one equation (check my arithmetic)

OpenStudy (anonymous):

The statement \[\frac{P(x)}{x+2}=4x^2+\frac{-ax-b}{x+2}\] comes from polynomial division. The fraction part is called the remainder. You cannot have a remainder in polynomial division if what you are dividing by is a factor of your polynomial. Because \(x+2\) is a factor, \(x+2\) should divide \(-ax-b\). Therefore, we say. \[-ax-b=x+2.\] Now, the only way that \(-ax-b\) is equal to \(x+2\) is if \(a=-1\) and \(b=-2.\) You can see this as follows \[-ax-b=-(-1)x-(-2)=x+2.\]

OpenStudy (anonymous):

replace \(x\) by 2 get \[4\times 2^3+8\times 2^2-2a-b=0\] \[64-2a-b=0\] \[-2a-b=-64\] \[2a+b=64\]

OpenStudy (anonymous):

solve the equations \[2a+b=64\] \[a+b=27\] and if my arithmetic is correct this should work quickly

OpenStudy (anonymous):

@satellite73, that is giving a different answer.. Hmm.

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

@Limitless did i mess up or is this arithmetic correct?

OpenStudy (anonymous):

I believe my answer is correct because the division works out in both cases. You must have messed up the arithmetic.

OpenStudy (anonymous):

it is late and hard for me to see the question and answer at the same time but i think this is the easiest method. you have two equations in two unknowns, should work quickly

OpenStudy (anonymous):

yeah i screwed up and used 2 instead of \(-2\) lets to it right

OpenStudy (anonymous):

@satellite73, I was wrong, actually. I messed up signs.

OpenStudy (anonymous):

@Limitless and @satellite73 so how do I find a and b your way?

OpenStudy (anonymous):

\[p(1)=-15=8+4-a-b\implies a+b=27\] \[p(-2)=4(-2)^3+8(-2)^2-2a-b=0\] \[-32+32-2a-b=0\] \[-2a-b=0\] \[2a+b=0\] we get two equations \[a+b=27\] \[2a+b=0\]

OpenStudy (anonymous):

subtract first equation from second get \(a=-27\) and so \(b=54\)

OpenStudy (anonymous):

That works, Sat. My method was incorrect.

OpenStudy (anonymous):

i didn't check to see if it works, but recall that \(a=-27\) so \(-a=27\) and further \(b=54\) so \(-b=-54\) and your polynomial should be \[p(x)=4x^3+8x^2+27x-54\] now we can check it

OpenStudy (anonymous):

Actually.. no, it doesn't. \(x+2\) does not divide \(P(x)\). http://www.wolframalpha.com/input/?i=%284x%5E3%2B8x%5E2%2B27x-54%29%2F%28x%2B2%29

OpenStudy (anonymous):

@purplec16, are you 100% sure there was not a typo? My method works if the remainder is \(15\), not \(-15.\)

OpenStudy (anonymous):

no there is not a typo... to my knowledge...

OpenStudy (anonymous):

But I'll ask...

OpenStudy (anonymous):

ok lets start from the top, but this time maybe we can write it as \[p(x)=4x^3+8x^2+ax+b\] to get rid of the annoying minus signs. obviously we just want \(a\) and \(b\)

OpenStudy (anonymous):

now you are told that \(x+2\) divides \(p(x)\)which means \(p(-2)=0\) so we replace \(x\) by -2, set the result equal zero and solve \[p(-2)=4(-2)^3+8(-2)^2+2a+b=0\] \[2a+b=0\]

OpenStudy (anonymous):

Okay.

OpenStudy (anonymous):

and you are told than when you divide by \(x-1\) you get \(-15\) which makes \(p(1)=-15\) we get \[4+8+a+b=-15\] \[12+a+b=-15\] \[a+b=-27\]

OpenStudy (anonymous):

two equations are \[2a+b=0\] \[a+b=-27\] subtract and get \[a=27, b=-54\] did i mess up yet?

OpenStudy (anonymous):

Well...before it was a=-27 and b=54

OpenStudy (anonymous):

oh i am an IDIOT!

OpenStudy (anonymous):

No you're not...

OpenStudy (anonymous):

\[\frac{P(x)}{x+2}=G(x)=4x^2+a+\frac{b-2a}{x+2}\] \[\frac{P(x)}{x-1}=H(x)=4x^2+12+a\frac{b-a+12}{x-1}\] \[b-2a=0\] \[b-a+12=-15\] \[b-a=-27\] ...

OpenStudy (anonymous):

\[-2a+b=0\]!!!

OpenStudy (anonymous):

ok i forgot that \(x=-2\) as the coefficient on the \(a\) term. god help me

OpenStudy (anonymous):

Aha! Yes...

OpenStudy (anonymous):

\[b-2-(b-a)=27 \Rightarrow -2+a=27 \Rightarrow a=29\] \[b-29=-27 \Rightarrow b=3\] Correct?

OpenStudy (anonymous):

It is really time for bed, smh

OpenStudy (anonymous):

\[b=2\]!!!

OpenStudy (anonymous):

Satellite, I cannot add. Do not worry.

OpenStudy (anonymous):

I really... don't understand how you came up with that @Limitless I am really new to this formula...

OpenStudy (anonymous):

And I forgot how to solve simultaneous equations..., smh

OpenStudy (anonymous):

@purplec16, we are both failing. This is us trying to work it out before explaining it.

OpenStudy (anonymous):

\(-2a+b=0\) \[a+b=-27\] subtract and get \(-3a=27\) so \(a=-9\) and \(b=18\) lets try that one

OpenStudy (anonymous):

Yeah, it should've been \(a+b+12=-15\).

OpenStudy (anonymous):

b=-18 you mean?

OpenStudy (anonymous):

yeash

OpenStudy (anonymous):

where'd the 12 come from? @Limitless

OpenStudy (anonymous):

Up above, polynomial division.

OpenStudy (anonymous):

And how do I factor the polynomial completely...

OpenStudy (anonymous):

We have finally accomplished this ;_;

OpenStudy (anonymous):

well you know it factors as \((x+2)(\text{something})\) i know right? arithmetic hell

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

you can find the "something" by polynomial division synthetic division thinking your choice although at this point thinking doesn't seem to be the best idea

OpenStudy (anonymous):

i would recommend synthetic division

OpenStudy (anonymous):

Thinking has demonstrated its own failings today.

OpenStudy (anonymous):

And what does the wolframalpha link show me?

OpenStudy (anonymous):

Okay I'll use synthetic

OpenStudy (anonymous):

but you know \[4x^2+8x^2-9x-18=(x+2)(ax^2+bx+c)\] and clearly \(a=4\) and \(c=-9\)

OpenStudy (anonymous):

so you only need to find \(b\) which you can do by thinking or rather should be able to

OpenStudy (anonymous):

yeah thinking did not serve me that well either

OpenStudy (anonymous):

Isn't it 4x^3+8x^2+9x+18?

OpenStudy (anonymous):

no i rewrote at \[4x^3+8x^2+ac+b\] because the minus signs were unnecessary and getting on my nerves

OpenStudy (anonymous):

it is \[4x^3+8x^2-9x-18=(x+2)(4x^2-9)=(x+2)(2x-3)(2x+3)\] it turns out \(b=0\) i this case

OpenStudy (anonymous):

Lol... okay... Thanks so much both @Limitless and @satellite73 idk who to give the medal to... :\ but you were both of such great help... I hope this is right, it sure looks so :)). Goodnight :).

OpenStudy (anonymous):

And I got that @satellite73

OpenStudy (anonymous):

lol we gave one to each other! so no need

OpenStudy (anonymous):

It is right, don't worry. You are welcome. And yeah, lol. We're sharing on this question!

OpenStudy (anonymous):

Okay :). Lol, thanks because I would have felt really bad if I only gave one and you both were equally helpful. Aww, that's so sweet. Thanks again!!!